Anns Antiques is estimating the following sales figures for the next four months: Cash sales Credit sales Total sales January | $1,000 20,000 21,000 February $1,300 $22,900 $24200 March 1.700 S23,500 $25,200 April | $2,100 | S24.000 25,100 Ann has the following collection pattern on credit sales: 15% in month of sale: 65% in month after sale: 20% in second month after sale. What are Anns budgeted cash receipts for March? $24,745 $24,110 $23,820 $22,665 QUESTION 14 Doras budgeted production figures (in units) for dresses for the next three months are shown below: Budgeted Production 5,000 June July 5,200 4,900 Dora uses 2.5 yards of fabric per dress and pays $3.00 per yard. Dora likes to have 30% of the next months fabric needs in ending inventory. what is the expected cost of fabric to be purchased in June? $50,025 $12,775 $16,675 $38,325
Answer #1

Question - 1

Budgeted receipt for march = Cash sales of march + 15%( march credit sales ) + 65% (Feb credit sales) + 20% ( Jan credit sales ) = 1700 + 0.15*23500 + 0.65 * 22900 + 0.20 * 20000 = 24110.............option - B

Question - 2

Purchase = [5200 + 30% (4900) - 30%(5200) ] * 2.5 = 4770 * 2.5 = 12775

Purchase in dollars = 12775 * 3 = $ 38325.............Option - D